Two numbers are such that the ratio between them is 3:5. If each is increased by 10, the ratio between the new numbers so formed is 5:7. Find the original numbers.

Answer by fcabanski(1391)

Equation 1: x/y = 3/5


5x = 3y


x = 3y/5


Equation 2:(x+10)/(y+10) = 5/7


7(x+10) =5(y+10)


7x + 70 = 5y + 50


Substitute in the value of x from equation 1.


7(3y/5) + 70 = 5y + 50


21y/5 + 70 = 5y + 50


20 = 5y - 21y/5


20 = 25y/5 - 21y/5 = 4y/5


100 = 4y


25 = y


Substitute that back into equation 1: x = 3*25/5 = 75/5 = 15